OpenStudy (anonymous):

Solve by the linear combination method (with or without multiplication). x + y = 50 0.05x + 0.06y = 2.8 A. (40, 60) B. (20, 30) C. (20, –50) D. (10, 30)

OpenStudy (anonymous):

@DSS I finally get the concept but now I just need someone who knows it better to double check because I can only get 2 wrong

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx again

OpenStudy (anonymous):

x + y = 50 I'd solve for x and use substitution. y = 50-x 0.05x + 0.06y = 2.8 0.05x +0.06(50-x)=2.8 x = 20 y = 30

OpenStudy (anonymous):

thx :D